OpenStudy (anonymous):

What is the ordered pair that is the solution to these equations. 2x + 5y = -9 3x - 4y = -2

OpenStudy (anonymous):

Multiply the first equation by 4 and the second by 5 8 x + 20 y = -36 15 x -20 y = -10 Add them up to find x, then find y.

OpenStudy (anonymous):

When you add then up, you get 23 x = - 46 and hence x=-2 replace in the first equation and you get -16 + 20 y = -36 20 y = -20 y=-1

OpenStudy (anonymous):

The final answer is x=-2 and y=-1